OpenStudy (anonymous):

SO M AND N UNDER =P-6 FOR N

hartnn (hartnn):

the question is \(\huge M=\frac{N}{P-6}\quad ?\)

hartnn (hartnn):

or \(\frac{M}{N}=P-6 \: ?\)

OpenStudy (anonymous):

YEA THAT ONE

hartnn (hartnn):

so 1. Multiply by N on both sides 2. and then divide by P-6 on both sides.

hartnn (hartnn):

u need help with that ^^ ?

OpenStudy (anonymous):

CAN YIU JUST DRAW THEM ALL

hartnn (hartnn):

can u just try the 1st part for me ? 1. Multiply by N on both sides

OpenStudy (anonymous):

I DID HATNM

OpenStudy (anonymous):

CON I TRIE IS IT M-P-6*N

hartnn (hartnn):

sorry, i kinda lost connection... so did u get M=N(P-6) ?

OpenStudy (anonymous):

ALLILTTLE DIFFRENT

OpenStudy (anonymous):

AND THEN

hartnn (hartnn):

now dividing (P-6) on both sides, u directly get N as N=\(\frac{M}{P-6}\)

OpenStudy (anonymous):

THATS THE ANSWER AND THEN SOLVEX-2 AND UNDER Y=Z FOR Y

hartnn (hartnn):

1. Multiply Y on both sides...

OpenStudy (anonymous):

SO X-2=YZ

hartnn (hartnn):

so u get X-2=YZ now divide Z on both sides...

OpenStudy (anonymous):

THATS WERE I GET STUCK NO YOU HAVE TO DRAW IT

hartnn (hartnn):